Poof:

2. Your bank account has in it. You deposit $5 per day. (a) How much money is in your account after 5 days? Show your work. (b) After you have made all of your deposits, you withdraw $2 a day until your account balance is $0. How many days will you withdraw $2? Show your work. Answer:

Poof:

your bank account has -15$

Poof:

Wow

AngeI:

5 per day for 5 days

AngeI:

5*5=25

AngeI:

Then you go down by 2's

Poof:

Mk

AngeI:

How many times can you do that before it hits 0

Poof:

Im horrible at this idek

AngeI:

In this case the easiest way to do it is just 25 divided by 2. Which will get you 12. You do have one more left over so you have to do another day because it is not at zero yet. So it would be 13 days for that. Unless you wants a decimal but I doubt that because it says 2 dollars every day and you would most likely withdraw it at the same time

Poof:

Wait if a have -15 dollars in my bank account and i desposit 5 per day how much is in my bank account after 5 days?

AngeI:

I suggest you delete that most recent part

AngeI:

I didnt see the 15 dollars thing

AngeI:

25-15=10

AngeI:

10/2=5

Poof:

okie

AngeI:

You should delete that file off of this post though

Poof:

okay

Poof:

There